OpenStudy (crashonce):

Hey guise FAN AND MEDAL!!!!!!! tanA+cosAsecBsecC=tanB+cosBsecAsecC and that it is also equal to tanC+cosCsecAsecB thanks

OpenStudy (crashonce):

@ikram002p

OpenStudy (crashonce):

@mathmate

OpenStudy (mathmate):

let me see...

OpenStudy (mathmate):

I would try the classic!

OpenStudy (crashonce):

what would that be? sorry im just a beginner

OpenStudy (mathmate):

not so sure any more about the classic.... decompose everything into sin and cos

OpenStudy (crashonce):

ok ill try that gimme a sec

OpenStudy (mathmate):

except that here we have three different angles, A, B and C, so they don't cancel that fast.

OpenStudy (mathmate):

You probably know something about what's relating A, B and C. Would A+B+C=180 degrees or something like that?

OpenStudy (crashonce):

wait sorry it says A+B+C=90

OpenStudy (mathmate):

By the symmetry of it, it sounds more like A=B=C.

OpenStudy (crashonce):

it says A+B+C=90 lol

OpenStudy (mathmate):

ok, at least there's something to go on! Have you done addition formulae?

OpenStudy (crashonce):

ya

OpenStudy (mathmate):

Wow, that's would be a big help. How do you think it could help?

OpenStudy (crashonce):

wait by addition formula do you mean cosx+cosy etc?

OpenStudy (mathmate):

more like tan(a+b)=(tan(a)+tan(b))/(1-tan(a)tan(b)) or something similar.

OpenStudy (crashonce):

yea i've got that but not sure how it helps

OpenStudy (mathmate):

Well, you could try eliminating A using A=90-B-C, and using addition formula to expand it, and hopefully something will cancel.

OpenStudy (crashonce):

but how is that meant to be an expansion?

OpenStudy (crashonce):

\[\cot(A+B) + (\sin(B+C)) \div cosBcosC\] thats what i have

OpenStudy (mathmate):

something like that, but we have to be careful which ones to choose.

OpenStudy (mathmate):

Perhaps we need to eliminate one of the variables A, B or C so we end up with only two variables.

OpenStudy (crashonce):

uh a few steps in, we get tan b and cosBsinCsecC but the expansion of the tangent value sorta messed it up

OpenStudy (mathmate):

I am still bugged by the symmetry of the three expression, which seems to tell me that I am missing something.

OpenStudy (crashonce):

yea idk either :(

OpenStudy (mathmate):

I will work on this on my own. In the mean time you can try gurus like @ganeshie8. Sorry! :(

OpenStudy (crashonce):

thats ok

OpenStudy (crashonce):

@ganeshie8 any ideas?

OpenStudy (crashonce):

@No.name idea?

OpenStudy (anonymous):

nopes sry

OpenStudy (crashonce):

thats ok

OpenStudy (crashonce):

@SolomonZelman

OpenStudy (crashonce):

anyone?

OpenStudy (crashonce):

@Purplerainbowcherry

OpenStudy (crashonce):

@Mimi_x3

OpenStudy (crashonce):

halp pls

OpenStudy (anonymous):

sorry :( I literally suck at these kind of questions :/

OpenStudy (mathmate):

@CrashOnce I cooked up something for you! To show: \(\large tanA+cosA secB secC = tanB+cosB seC secA = tanC+cosC secA secB\) To do this, it would be less confusing to prove two simple relations ahead of time to simplify the proof. These two proofs (called lemmas) require the use of addition/subtraction trigonometric formulas: \(\large cos(B+C) = cosB cosC - sinB sinC\) from which \(\large cosB cosC = cos(B+C) + sinB sinC\)....(1) From A+B+C=90, and \(\large cos(X-Y) = cosX cosY + sinX sinY\) We have \(\large cos(B+C)=cos(90-A) = cos90 cosA + sin90 sinA\) since cos90=0, sin90=1 \(\large cos(B+C)= sinA\) ....(2) Starting with the first expression: \(\large tanA + cosA secB secC\) We reduce to sin and cos to help simplify things \(\large \frac{sinA}{cosA} + \frac{cosA}{cosB cosC}\) join expressions using common denominator: \(\large \frac{sinA cosB cosC + cos^2A}{cosA cosB cosC}\) Apply relation (1) above to cosB cosC: \(\large \frac{sinA(cos(B+C) + sinB sinC) ) + cos^2A}{cosA cosB cosC}\) Apply relation (2) above to cos(B+C): \(\large \frac{sinA(sinA + sinB sinC) + cos^2A}{cosA cosB cosC}\) Expanding: \(\large \frac{sin^2A + sinA sinB sinC + cos^2A}{cosA cosB cosC}\) \(\large \frac{sin^2A + cos^2A + sinA sinB sinC}{cosA cosB cosC}\) Using \(sin^2+cos^A = 1\) \(\large \frac{1+sinA sinB sinC}{cosA cosB cosC}\)...(3) Hence \(\large tanA+cosA secB secC \equiv \frac{1+sinA sinB sinC}{cosA cosB cosC}\) ...(4) We can similarly show that \(\large tanB+cosB secC secA \equiv \frac{1+sinB sinC sinA}{cosB cosC cosA}\) ...(5) and \(\large tanC+cosC secA secB \equiv \frac{1+sinC sinA sinB}{cosC cosA cosB}\) ...(6) Since (4),(5),(6) all equal to the same expression symmetric in A,B and C, the given identity is proved.

OpenStudy (crashonce):

ah thanks @mathmate

OpenStudy (mathmate):

You're welcome! :)